OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 3 –2 4 Line 2 x y 0 8 2 6 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Do you want an explanation or the answer?

OpenStudy (anonymous):

@Directrix @Godlovesme

OpenStudy (anonymous):

would you like to expain it so i can "get it"

OpenStudy (anonymous):

your statement last time confused me so much

OpenStudy (anonymous):

much more than i was

OpenStudy (anonymous):

let's say we have two equations that represent lines: y=5x+2 y=2x+1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If they asked you how many solutions this has, could you do it?

OpenStudy (anonymous):

i would think 2

OpenStudy (anonymous):

A solution graphically means how many times the lines intersect.

OpenStudy (anonymous):

meaning how many times the lines meet right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):



OpenStudy (anonymous):

So if we wanted to know how many solutions (for example) y=5x+2 y=2x+1 These two equations have, we would say their y is the same, and simply equate them. 5x+2 = 2x+1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):